OpenStudy (anonymous):

A. find surface area of the cube. B. find the volume of the cube. one side is 4x-5

OpenStudy (anonymous):

surface area = 6s^2 where s is the side ... volume=s^3 i hope you can handle it now

OpenStudy (anonymous):

Let side of the cube 4x-5 = a So Surface Area = 6a^2 ==> 6(4x-5)^2 ==> 6(16x^2 - 40x + 25) ==> 96x^2 - 240x + 150 And Volume = a^3 ==> (4x - 5)^3 ==> 64x^3 - 240x^2 + 225x - 125

OpenStudy (anonymous):

is this a test ur takin or an assignment??

OpenStudy (anonymous):

assignment trying to help my son

OpenStudy (anonymous):

are those the final answers?